# ToTRL: Unlock LLM Tree-of-Thoughts Reasoning Potential through Puzzles Solving

### v1

@@ 1 Introduction
Large language models (LLMs) have recently demonstrated remarkable capabilities in tackling complex reasoning tasks, with many advanced LLMs [3, 9, 6, 8] commonly employing the chain-of-thought (CoT) [18] technique to generate explicit, step-by-step intermediate reasoning. This sequential reasoning process has enabled powerful inference in structured domains like mathematics and programming. For example, the GPT-o1 series [13] achieves improved inference performance with longer reasoning chains. Similarly, DeepSeek-R1 [9] attains notable results in complex reasoning via the emergent long CoT reasoning process, which is activated using reinforcement learning (RL) with a rule-based reward signal.
@@
However, the fundamental nature of CoT is a linear and single-path reasoning process. Although effective for certain problems, this sequential structure inherently limits its efficiency when addressing tasks that necessitate exploring and evaluating multiple potential solution trajectories. A critical issue arises when using RL with rule-based rewards to induce longer reasoning, the resulting outputs can become verbose and redundant [13, 9, 6]. This is because the underlying reasoning remains local, moving from one step to the next without a global perspective or an effective mechanism to evaluate the overall promise of a path or prune unpromising lines of thought. This contrasts with the efficient human cognitive strategy, which involves considering alternatives and focusing resources globally.
@@
In contrast, the tree-of-thoughts (ToT) method [22] offers a conceptually superior approach by explicitly modeling the reasoning process as an exploration across a tree structure of potential thoughts or states. The tree-based reasoning structure facilitates the parallel generation and evaluation of diverse reasoning branches, enabling the LLM to actively identify, evaluate, and prune unproductive thought paths. By maintaining a global view of the search space, ToT holds the potential to achieve higher performance and significantly reduce redundant exploration and associated token costs compared to the linear CoT reasoning process.
@@
Building upon the capabilities of long CoT [14], we introduce tree-of-thoughts RL (ToTRL), a novel on-policy RL framework using the rule-based reward. ToTRL is designed to guide LLMs in developing the parallel ToT strategy [22] based on the sequential CoT strategy [23]. Directly building ToT reasoning based on CoT within the reasoning mode is challenging due to the established habituation to the sequential CoT style. To address this, ToTRL employs a two-stage training strategy. Initially, the LLM is trained to perform ToT reasoning in a non-thinking mode, leveraging more moldable thinking patterns to activate ToT reasoning. Once the LLM has developed a degree of ToT reasoning ability in the non-reasoning mode, it undergoes further training in the reasoning mode. This second stage aims to enable the LLM to effectively utilize its newly acquired ToT capabilities during inference on complex tasks, building upon its existing CoT reasoning strengths.
@@
Moreover, we employ LLMs as players in puzzle games that require tree-based reasoning, providing challenging tasks for cultivating the ToT reasoning capability. These puzzle games are deliberately chosen for their intrinsic requirement of exploring interdependent choices and managing multiple constraints simultaneously, requiring a thought tree construction and exploration involving multiple concurrent hypotheses and future states. Our contributions are summarized as follows:
@@ 2 Tree-of-Thoughts RL
Building upon the capabilities of long CoT [14], we introduce tree-of-thoughts RL using the on-policy RL strategy (Section 2.1) with the rule-based reward (Section 2.2). Specifically, ToTRL employs a two-stage training strategy (Section 2.3) and serves LLMs as the puzzle game player for training (Section 2.4).
@@ 2.1 On-policy RL Algorithm
Test-time scaling introduces a significant paradigm shift for LLMs [13, 9]. This approach enables long CoT reasoning and fosters sophisticated reasoning behaviors, leading to superior performance on complex reasoning tasks. A key technique facilitating these advancements is rule-based RL, which elicits behaviors such as self-verification and iterative refinement.
In this paper, we employ REINFORCE [27, 10] as the on-policy RL algorithm for our proposed ToTRL method. Specifically, for each prompt $q$, we sample $n$ responses $\{o_{1},o_{2},\dots,o_{n}\}$ from the old policy $\pi_{\theta_{\text{old}}}$, where $n$ is the number of sampled trajectories (i.e., the rollout size per prompt). The policy model $\pi_{\theta}$ is then optimized by maximizing the following surrogate objective:
$\mathcal{J}(\theta)$ $=\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{n}\sim\pi_{\theta_{\text{old}}}(o|q)}$
$\frac{1}{n}\sum\limits_{i=1}^{n}\bigl{(}\min\bigl{(}\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}A_{i},\text{clip}\bigl{(}\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)},1-\epsilon,1+\epsilon\bigr{)}A_{i}\bigr{)}-\beta\cdot\mathbb{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\bigr{)},$ (1)
@@
Traditionally, RL algorithms incorporate a KL divergence penalty to regulate the divergence between the online policy model and the frozen reference model [24]. However, during training with ToTRL, the model distribution can diverge significantly from the initial model. The KL penalty term will restrict the exploration of model outputs. Consequently, we exclude the KL term by setting $\beta=0$.
@@ 2.2 Reward Modeling
$R(o|q;r_{\text{s}},r_{\text{p}})=\begin{cases}r_{\text{p}},&\text{if }\mathcal{F}(o)=0,\\ \mathcal{A}(o|x)\cdot r_{\text{s}}+(1-\mathcal{A}(o|x))\cdot r_{\text{p}},&\text{if }\mathcal{F}(o)=1.\end{cases}$ (6)
Alternatively, the total reward can be expressed in a mathematically equivalent and compact form as:
@@ 2.3 Multi-Stage ToTRL
Conventional CoT reasoning enforces a strictly linear, step-by-step reasoning process [18, 23]. This linearity can be restrictive when exploring multiple potential solution paths. Inspired by ToT [22], which structures problem-solving as a deliberate exploration of a thought tree, ToTRL adapts this concept to enhance LLM reasoning within an RL setting. The original ToT framework represents the reasoning process as a rooted tree, where each node corresponds to an intermediate thought $s$. From a state $s_{d-1}$ at depth $d-1$, a generator $G(\cdot)$ proposes potential next thoughts, forming a set of candidate states at depth $d$:
@@
In our ToTRL approach, instead of implementing an explicit external search algorithm, we develop a ToT guidance prompt to facilitate the parallel generation of ToT reasoning processes within each RL rollout. As depicted in Figure 1, LLMs are required to solve problems by employing ToT reasoning, ensuring each step is executed sequentially without omitting any intermediate stages. This prompt-driven process encourages the LLM to explore ToT reasoning trajectories autonomously within its own generation process. The output generated after ToT reasoning is subsequently evaluated using Equation 6, and the entire process is optimized via Equation 1.
However, existing reasoning LLMs are primarily accustomed to the sequential CoT reasoning style [14]. Consequently, integrating ToT reasoning directly into current CoT-based reasoning LLMs presents a significant challenge. To address this challenge, ToTRL employs a multi-stage ToT guidance strategy. Initially, as illustrated in Figure 1, the LLM undergoes training to perform ToT reasoning in a non-thinking mode. The non-reasoning mode is achieved by introducing blanks between reasoning tags (e.g., <think></think>), which compels the model to suspend its conventional reasoning processes. Once the LLM demonstrates an initial proficiency in ToT reasoning within the non-reasoning mode, it proceeds to further training in the reasoning mode. This subsequent stage aims to enable the LLM to develop its newly acquired ToT capabilities based on the established CoT reasoning strengths.
@@ 2.4 LLM as Puzzle Game Player
During the training process with our ToTRL, LLMs are employed as players in puzzle games. These games necessitate tree-based reasoning, thereby providing challenging tasks for cultivating the LLMs’ ToT reasoning capabilities. They are specifically selected due to their inherent demand for exploring interdependent choices and managing multiple concurrent constraints, which inherently requires the construction and exploration of a sophisticated thought tree involving numerous concurrent hypotheses and future states. Specifically, as shown in Figure 2, we leverage Sudoku and Alphametic puzzles during the training process with ToTRL for this purpose.
@@
By employing LLMs as players in these puzzle games, we provide challenging environments designed to foster their tree-based reasoning capabilities within complex constraint satisfaction problems.
@@ 3.1 Experiment Settings
Training Data. The ToTQwen3-8B model is trained using 1440 puzzle games, specifically designed to cultivate its ToT reasoning capabilities. We provide more details in Section B.1.
Implementation Details. In this paper, we train the ToTQwen3-8B model using our ToTRL framework based on the Qwen3-8B model [14]. The training process entailed using a constant learning rate schedule with a warm-up ratio of 0.01, and the AdamW [12] optimizer with a learning rate of $1\times 10^{-6}$, no weight decay, a batch size of 9, and a sequence length of 16384 tokens. The models underwent instruction tuning for one epoch using DeepSpeed-Zero stage2 with offload [15] on 4 A100 GPUs, each with 80G memory.
Baselines. We select the existing SOTA reasoning LLMs with similar parameters as baselines of our ToTQwen3-8B model, which include DeepSeek-R1-Distill-Qwen-7B [9], Llama-3.1-Nemotron-Nano-8B [4], GLM-4-Z1-9B-0414 [28], Phi-4 Reasoning [1], and Qwen3-8B [14]. Notably, Qwen3-8B and ToTQwen3-8B use the thinking mode for evaluation.
Evaluation Benchmarks. To comprehensively evaluate the proficiency of our ToTQwen3-8B model, we utilize a set of logic reasoning tasks categorized as either in-distribution or out-of-distribution (OOD). In-distribution tasks, included in the training data, comprise 6$\times$6 Sudoku and Alphametic puzzles. OOD tasks, which are not part of the training data, include 5$\times$5 Crossword [22], 9$\times$9 Sudoku [5], K&K puzzles [21], Poker 24 Game, and Make 24 puzzles. Additionally, we also leverage the widely adopted AIME 2024–2025 (American Invitational Mathematics Examination) and AMC 2023 (American Mathematics Competitions) benchmarks, both known for their rigorous and diverse mathematical problems. Specifically, for the evaluation of our ToTQwen3-8B, we employ the ToT prompt (Figure 1) for logic reasoning tasks while leveraging the CoT prompt for mathematical problems. For all evaluation tasks, we use accuracy as the performance metric. Notably, the accuracy is calculated on the number of correct answers for puzzles with multiple solutions.
Thinking Budget. Thinking budget facilitates control over the thinking process through manual interruption. Specifically, when the LLM’s thinking duration reaches a predefined threshold, the process is halted by inserting a stop instruction. Subsequently, the LLM generates a final response based on the partial reasoning accumulated. Notably, this budget was adjusted based on task complexity. A smaller budget of 8K is allocated for the simple K&K Puzzle, while a larger budget of 32K is used for the more complex 9$\times$9 Sudoku.
@@ 3.2 In-Distribution Logic Reasoning
As shown in Table 1, performance on these logic tasks varies across the evaluated models. Notably, the ToTQwen3-8B model demonstrates significantly superior performance in in-distribution tasks compared to the other evaluated models on both tasks. Specifically, ToTQwen3-8B achieves the highest success rate on the 6$\times$6 Sudoku task. On the Alphametic Puzzles, it consistently performs at a high level across puzzles with varying numbers of solutions.
@@
6$\times$6 Sudoku Alphametic Puzzle
@@
5$\times$5 Crossword 9$\times$9 Sudoku K&K Puzzle

### v2

@@ Abstract
[…]
ToTRL: Unlock LLM Tree-of-Thoughts Reasoning Potential through Puzzles Solving
[…]
@@ 1 Introduction
Large language models (LLMs) have recently demonstrated remarkable capabilities in tackling complex reasoning tasks, with many advanced LLMs (anthropic2025claude; guo2025deepseekr1; kimi2025k1.5; deepmind2025gemini) commonly employing the chain-of-thought (CoT) (wei2022cot) technique to generate explicit, step-by-step intermediate reasoning. This sequential reasoning process has enabled powerful inference in structured domains like mathematics and programming. For example, the GPT-o1 series (openai2024o1) achieves improved inference performance with longer reasoning chains. Similarly, DeepSeek-R1 (guo2025deepseekr1) attains notable results in complex reasoning via the emergent long CoT reasoning process, which is activated using reinforcement learning (RL) with a rule-based reward signal.
@@
However, the fundamental nature of CoT is a linear and single-path reasoning process. Although effective for certain problems, this sequential structure inherently limits its efficiency when addressing tasks that necessitate exploring and evaluating multiple potential solution trajectories. A critical issue arises when using RL with rule-based rewards to induce longer reasoning, the resulting outputs can become verbose and redundant (openai2024o1; guo2025deepseekr1; kimi2025k1.5). This is because the underlying reasoning remains local, moving from one step to the next without a global perspective or an effective mechanism to evaluate the overall promise of a path or prune unpromising lines of thought. This contrasts with the efficient human cognitive strategy, which involves considering alternatives and focusing resources globally.
@@
In contrast, the tree-of-thoughts (ToT) (yao2023tot) method offers a conceptually superior approach by explicitly modeling the reasoning process as an exploration across a tree structure of potential thoughts or states. The tree-based reasoning structure facilitates the parallel generation and evaluation of diverse reasoning branches, enabling the LLM to actively identify, evaluate, and prune unproductive thought paths. By maintaining a global view of the search space, ToT holds the potential to achieve higher performance and significantly reduce redundant exploration and associated token costs compared to the linear CoT reasoning process.
@@
Building upon the capabilities of long CoT (qwen2025qwen3), we introduce tree-of-thoughts RL (ToTRL), a novel on-policy RL framework using the rule-based reward. ToTRL is designed to guide LLMs in developing the parallel ToT strategy (yao2023tot) based on the sequential CoT strategy (yeo2025longcot). Directly building ToT reasoning based on CoT within the reasoning mode is challenging due to the established habituation to the sequential CoT style. To address this, ToTRL employs a two-stage training strategy. Initially, the LLM is trained to perform ToT reasoning in a non-thinking mode, leveraging more moldable thinking patterns to activate ToT reasoning. Once the LLM has developed a degree of ToT reasoning ability in the non-reasoning mode, it undergoes further training in the reasoning mode. This second stage aims to enable the LLM to effectively utilize its newly acquired ToT capabilities during inference on complex tasks, building upon its existing CoT reasoning strengths.
@@
Moreover, we employ LLMs as players in puzzle games that require tree-based reasoning, providing challenging tasks for cultivating the ToT reasoning capability. These puzzle games are deliberately chosen for their intrinsic requirement of exploring interdependent choices and managing multiple constraints simultaneously, requiring a thought tree construction and exploration involving multiple concurrent hypotheses and future states.
Our contributions are summarized as follows:
@@ 2 Tree-of-Thoughts RL
Building upon the capabilities of long CoT (qwen2025qwen3), we introduce tree-of-thoughts RL using the on-policy RL strategy (Section˜2.1) with the rule-based reward (Section˜2.2). Specifically, ToTRL employs a two-stage training strategy (Section˜2.3) and serves LLMs as the puzzle game player for training (Section˜2.4).
@@ 2.1 On-policy RL Algorithm
Test-time scaling introduces a significant paradigm shift for LLMs (openai2024o1; guo2025deepseekr1). This approach enables long CoT reasoning and fosters sophisticated reasoning behaviors, leading to superior performance on complex reasoning tasks. A key technique facilitating these advancements is rule-based RL, which elicits behaviors such as self-verification and iterative refinement.
In this paper, we employ the on-policy RL algorithm (schulman2017ppo; zhang2021reinfore; hu2025reinforce++; ahmadian2024rloo) for our proposed ToTRL method. Specifically, for each prompt $q$, we sample $n$ responses $\{o_{1},o_{2},\dots,o_{n}\}$ from the old policy $\pi_{\theta_{\text{old}}}$, where $n$ is the number of sampled trajectories (i.e., the rollout size per prompt). The policy model $\pi_{\theta}$ is then optimized by maximizing the following surrogate objective:
$q\sim$ $\mathcal{D},\;\{o_{i}\}_{i=1}^{n}\sim\pi_{\theta_{\text{old}}}(o|q),\;r_{i}=\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)},$
$L_{i}$ $=\min\bigl(r_{i}A_{i},\text{clip}(r_{i},1-\epsilon,1+\epsilon)A_{i}\bigr),$
$\mathcal{J}(\theta)$ $=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\left(L_{i}-\beta\cdot\mathbb{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\right)\right].$ (1)
@@
Traditionally, RL algorithms incorporate a KL divergence penalty to regulate the divergence between the online policy model and the frozen reference model (yu2025dapo). However, during training with ToTRL, the model distribution can diverge significantly from the initial model. The KL penalty term will restrict the exploration of model outputs. Consequently, we exclude the KL term by setting $\beta=0$.
@@ 2.3 Multi-Stage ToTRL
Conventional CoT reasoning enforces a strictly linear, step-by-step reasoning process (wei2022cot; yeo2025longcot). This linearity can be restrictive when exploring multiple potential solution paths. Inspired by ToT (yao2023tot), which structures problem-solving as a deliberate exploration of a thought tree, ToTRL adapts this concept to enhance LLM reasoning within an RL setting. The original ToT framework represents the reasoning process as a rooted tree, where each node corresponds to an intermediate thought $s$. From a state $s_{d-1}$ at depth d1d-1, a generator $G(\cdot)$ proposes potential next thoughts, forming a set of candidate states at depth $d$:
@@
In our ToTRL approach, instead of implementing an explicit external search algorithm, we develop a ToT guidance prompt to facilitate the parallel generation of ToT reasoning processes within each RL rollout. As depicted in Figure˜1, LLMs are required to solve problems by employing ToT reasoning, ensuring each step is executed sequentially without omitting any intermediate stages. This prompt-driven process encourages the LLM to explore ToT reasoning trajectories autonomously within its own generation process. The output generated after ToT reasoning is subsequently evaluated using Equation˜6, and the entire process is optimized via Equation˜1.
However, existing reasoning LLMs are primarily accustomed to the sequential CoT reasoning style (qwen2025qwen3). Consequently, integrating ToT reasoning directly into current CoT-based reasoning LLMs presents a significant challenge. To address this challenge, ToTRL employs a multi-stage ToT guidance strategy. Initially, as illustrated in Figure˜1, the LLM undergoes training to perform ToT reasoning in a non-thinking mode. The non-reasoning mode is achieved by introducing blanks between reasoning tags, which compels the model to suspend its conventional reasoning processes. Once the LLM demonstrates an initial proficiency in ToT reasoning within the non-reasoning mode, it proceeds to further training in the reasoning mode. This subsequent stage aims to enable the LLM to develop its newly acquired ToT capabilities based on the established CoT reasoning strengths.
@@ 2.4 LLM as Puzzle Game Player
During the training process with our ToTRL, LLMs are employed as players in puzzle games. These games necessitate tree-based reasoning, thereby providing challenging tasks for cultivating the LLMs’ ToT reasoning capabilities. They are specifically selected due to their inherent demand for exploring interdependent choices and managing multiple concurrent constraints, which inherently requires the construction and exploration of a sophisticated thought tree involving numerous concurrent hypotheses and future states. Specifically, as shown in Figure˜2, we leverage Sudoku and Alphametic puzzles during the training process with ToTRL for this purpose.
@@
Employing LLMs as players in puzzle games, we provide challenging environments designed to foster their tree-based reasoning capabilities within complex constraint satisfaction problems.
@@ 3.1 Experiment Settings
Training Data. The ToTQwen3-8B model is trained using 1440 puzzle games, specifically designed to cultivate its ToT reasoning capabilities. We provide more details in Section˜B.1.
Implementation Details. We train the ToTQwen3-8B model using our ToTRL framework, initializing it from the Qwen3-8B model (qwen2025qwen3). The training process utilized the AdamW optimizer (loshchilov2017adamw) with a constant learning rate of $1\times 10^{-6}$, incorporating a warm-up ratio of 0.01. We employed a batch size of 9, a rollout size of 16, and a maximum sequence length of 16384 tokens, with no weight decay applied. The model undergoes full-parameter fine-tuning for one epoch using DeepSpeed-Zero stage 2 with CPU offload (rajbhandari2020zero).
Baselines. We select the existing SOTA reasoning LLMs with similar parameters as baselines of our ToTQwen3-8B model, which include DeepSeek-R1-Distill-Qwen-7B (guo2025deepseekr1), Llama-3.1-Nemotron-Nano-8B (bercovich2025llama_nemotron), GLM-4-Z1-9B-0414 (zhipu2025glm0414), Phi-4 Reasoning (abdin2025phi4reasoning), and Qwen3-8B (qwen2025qwen3). Notably, Qwen3-8B and ToTQwen3-8B use the thinking mode for evaluation.
Evaluation Benchmarks. To comprehensively evaluate the proficiency of our ToTQwen3-8B model, we utilize a set of logic reasoning tasks categorized as either in-distribution or out-of-distribution (OOD). In-distribution tasks, included in the training data, comprise $\times6$ Sudoku and Alphametic puzzles. OOD tasks, which are not part of the training data, include $\times5$ Crossword (yao2023tot), $\times9$ Sudoku (kaggle9dataset), K&K puzzles (xie2025logicrl), Poker 24 Game, and Make 24 puzzles. Additionally, we also leverage the widely adopted AIME 2024–2025 (American Invitational Mathematics Examination) and AMC 2023 (American Mathematics Competitions) benchmarks, both known for their rigorous and diverse mathematical problems. Specifically, for the evaluation of our ToTQwen3-8B, we employ the ToT prompt (Figure˜1) for logic reasoning tasks while leveraging the CoT prompt for mathematical problems. For all evaluation tasks, we use accuracy as the performance metric. Notably, the accuracy is calculated on the number of correct answers for puzzles with multiple solutions.
Thinking Budget. Thinking budget facilitates control over the thinking process through manual interruption. Specifically, when the LLM’s thinking duration reaches a predefined threshold, the process is halted by inserting a stop instruction. Subsequently, the LLM generates a final response based on the partial reasoning accumulated. Notably, this budget was adjusted based on task complexity. A smaller budget of 8K is allocated for the simple K&K Puzzle, while a larger budget of 32K is used for the more complex $\times9$ Sudoku. We provide more details about thinking budget in Section˜B.3.
@@ 3.2 In-Distribution Logic Reasoning
As shown in Table˜1, performance on these logic tasks varies across the evaluated models. Notably, the ToTQwen3-8B model demonstrates significantly superior performance in in-distribution tasks compared to the other evaluated models on both tasks. Specifically, ToTQwen3-8B achieves the highest success rate on the $\times6$ Sudoku task. On the Alphametic Puzzles, it consistently performs at a high level across puzzles with varying numbers of solutions.
@@
$\times6$ Sudoku Alphametic Puzzle
@@
$\times5$ Crossword $\times9$ Sudoku K&K Puzzle
@@
3.3 Out-of-Distribution Logic Reasoning
Besides in-distribution tasks, we also evaluate our ToTQwen3-8B model on OOD logic reasoning tasks. As shown in Table˜2 and Table˜3, many models exhibit very low scores (close to 0), indicating a significant struggle with these logic reasoning tasks.
Logic Reasoning with Unique Solution. Table˜2 presents performance on OOD logic reasoning tasks, all of which have a single correct solution. It is important to note that for tasks with unique solutions, some models might occasionally arrive at the correct answer through guessing rather than a complete logical derivation process. Our ToTQwen3-8B consistently achieves the highest scores across all three unique solution tasks evaluated, 0.508 on $\times5$ Crossword, 0.260 on $\times9$ Sudoku, and 0.986 on K&K Puzzle. This consistent, high performance on diverse OOD tasks provides strong evidence that our ToTQwen3-8B model effectively leverages the underlying logical constraints and employs a well-developed derivation process, rather than relying on chance.
Logic Reasoning with Multiple Solutions. Table˜3 presents performance on more OOD logic reasoning tasks, including Poker 24 Game and Make 24 Puzzle, which admit multiple valid solutions. The primary challenge lies not only in identifying a valid solution but also potentially in finding multiple distinct solutions or performing robustly across puzzles with varying numbers of possible solutions. This necessitates a reasoning process capable of exploring diverse successful paths rather than merely converging on a single outcome. As shown in Table˜3, ToTQwen3-8B again outperforms all other models on both the Poker 24 Game and Make 24 Puzzle, demonstrating superior performance across puzzles with varying numbers of solutions. This advantage is most pronounced on puzzles with many solutions, where the capacity to explore multiple valid reasoning paths is essential. The ToT strategy inherently explores a tree of possibilities, branching at critical decision points. This systematic exploration of multiple potential continuations provides a global perspective, making it inherently well-suited for tasks admitting diverse valid solutions. Unlike long CoT reasoning strategy that commits to a single path, ToT can explore parallel branches, significantly increasing the probability of discovering multiple distinct solutions when they exist.
In summary, from an OOD perspective, the results presented in Table˜2 and Table˜3 strongly indicate that ToTQwen3-8B possesses superior generalization capabilities on these logic reasoning tasks. Its robust performance on both unique and multiple-solution puzzles suggests that the ToT reasoning structure confers a significant advantage in addressing unfamiliar logic challenges by facilitating a more comprehensive exploration of reasoning possibilities, a capability that is particularly valuable for tasks requiring the discovery of multiple valid solutions.
Figure 3: Illustration of test time scaling on logic reasoning tasks.
3.4 Out-of-Distribution Mathematical Tasks
We further investigate whether our ToTRL can enhance complex reasoning abilities can transfer to a highly challenging mathematical reasoning scenario. According to Table˜4, ToTQwen3-8B demonstrates strong performance on these OOD mathematical reasoning tasks. It achieves the highest average score of 0.750 across the three benchmarks. Specifically, ToTQwen3-8B performs comparably well on AIME 2024 with a score of 0.667, matching the performance of GLM-4-Z1-9B-0414 (zhipu2025glm0414) and Phi-4 Reasoning (abdin2025phi4reasoning) with more model parameters. On AIME 2025, ToTQwen3-8B leads with a score of 0.633, outperforming all other listed models. For the AMC 2023 benchmark, ToTQwen3-8B achieves a score of 0.950, which is tied with GLM-4-Z1-9B-0414 (abdin2025phi4reasoning) for the highest performance.
The ToT training process encourages the LLM to generate more diverse, parallel, and structured intermediate reasoning steps, with each parallel branch employing the CoT format. This approach ultimately improves the quality of the sequential reasoning steps generated within the CoT prompt, which is crucial for tackling complex OOD problems. The strong performance observed on OOD tasks indicates that our ToTRL framework effectively enhances the LLM’s reasoning capabilities. Importantly, these improved abilities generalize well to novel and challenging mathematical problems beyond the training distribution, suggesting that ToTRL successfully cultivates a more robust and transferable form of reasoning.
3.5 Test Time Scaling
Test-time scaling (yu2025z1) is an innovative approach in language modeling that leverages additional computational resources during the testing phase to enhance performance. This method has shown significant promise in various domains, including language modeling and code generation (openai2024o1; guo2025deepseekr1; kimi2025k1.5; deepmind2025gemini). We investigate the test time scaling of our proposed ToTQwen3-8B model against a baseline Qwen3-8B (qwen2025qwen3) model using different thinking budgets across six logic reasoning tasks.
As illustrated in Figure˜3, the performance of both ToTQwen3-8B and Qwen3-8B generally improves with an increased thinking budget across all tasks. This indicates that allowing more computational resources for intermediate thinking steps leads to better reasoning outcomes. Importantly, Figure˜3 demonstrates the efficiency of the ToTQwen3-8B approach, which leverages the ToT reasoning strategy. ToTQwen3-8B consistently outperforms Qwen3-8B across various thinking budgets. Notably, ToTQwen3-8B is often able to reach a higher average score with a smaller thinking budget compared to Qwen3-8B. This suggests that the structured tree-of-thoughts reasoning employed by ToTQwen3-8B allows it to explore the solution space more effectively and efficiently, requiring fewer tokens (and thus less computational cost and time) to achieve superior or comparable performance. This efficiency is a key advantage, making ToTQwen3-8B a more practical solution for logic reasoning tasks under computational constraints.